How 0+0=1

How 0+0=1

Grade:7

3 Answers

shekar
10 years ago
it is so silly......can u pls reveal the answer.
kumar
10 years ago
0+0=1 it can be written as (2-2)+(2-2)=1 2-2+2-2=1 4-4=1 4=1+4 5=4 5-4=0 1=0+0 hence it is proved
PERUMAL
10 years ago
0+0=1 this, can be written as, (2-2+(2-2)=1 (4-4)=1 4=1+4 4=5 5-4=0 1=0+0 hense, this problem prove.